OpenStudy (anonymous):

Add. Simplify by collecting like radical terms if possible, assuming all expressions under radicands express nonnegative number: sqrt{5a}+3sqrt{125a^{3}}

OpenStudy (anonymous):

sqrt{5a}+3sqrt{125a^{3}}

OpenStudy (anonymous):

grr it wont post it right XD \[\sqrt{5a}+3\sqrt{125a^{3}} \]

OpenStudy (anonymous):

\[125=5^3\] so you get \[\sqrt{5a}+15a\sqrt{5a}\] to combine write \[(1+15a)\sqrt{5a}\]

OpenStudy (anonymous):

I got \[15\sqrt{5a}\] but I have and icky feeling Im wrong

OpenStudy (anonymous):

you have to take an a out as well

OpenStudy (anonymous):

oooh so very close! thanks again satellite!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

while I have you here can I ask another? you seem to be comfortable with these